OpenStudy (anonymous):

A sample of 20.0 moles of a monatomic ideal gas (γ = 1.67) undergoes an adiabatic process. The initial pressure is 400kPa and the initial temperature is 450K. The final temperature of the gas is 320K. What is the final volume of the gas? Let the ideal-gas constant R = 8.314 J/(mol • K).

OpenStudy (aaronq):

\[\frac{ T _{2} }{ T _{1} }=\left( \frac{ V _{1} }{ V _{2} } \right)^{\gamma-1}\] you can find the derivation of this in your book, or the internet (google "adiabatic compression")

OpenStudy (aaronq):

ps find the initial volume first

OpenStudy (anonymous):

I did: PV = nRT P = 400 kPa n = 20.0 T = 450 K V = nRT/P = 20 * 8.314 * 450 / 400 = 187 L V1/T1 = V2/T2 and I'm solving for V2 V2 = V1 (T2/T1) = 187 x 320 / 450 = 133 L My answer doesn't match the choices. I don't know what I did wrong.): the choices are: A. 230 L B. 350 L C. 270 L D. 190 L E. 310 L

OpenStudy (aaronq):

you didn't take into account the exponent, gamma-1 on the (V1/V2)

OpenStudy (anonymous):

Does gamma-1= -0.422784?

OpenStudy (anonymous):

Is this gamma the same thing as euler's constant?

OpenStudy (aaronq):

nope, gamma (the heat capacity ratio) was given in the question. http://en.wikipedia.org/wiki/Heat_capacity_ratio

OpenStudy (anonymous):

Ohhh, I didn't even notice the 1.67. Thank you for your help!

OpenStudy (aaronq):

no problem! read the question next time :P